please help me solve using subsititution method
7x+5y=10
x=31-4y
*solution is ordered pair*

Answer by rfer(16322)

7x+5y=10
x=31-4y
-------------
7(31-4y)+5y=10
217-28y+5y=10
-23y=-207
y=9
----------------
x=31-4(9)
x=31-36
x=-5
---------------
(-5,9)